OpenStudy (anonymous):

can someone help me integrate x/(2x-1)

OpenStudy (anonymous):

u must integrate by part. u = x dv = dx/(2x-1) if you have specific questions let me know

OpenStudy (tkhunny):

Or, long division has been working nicely for generations. \(\dfrac{x}{2x-1} = \dfrac{1}{2} + \dfrac{1/2}{2x-1}\)

OpenStudy (anonymous):

Or a substitution! Lots of ways to go about it. Let \(u=2x-1~~\Rightarrow~~\dfrac{1}{2}du=dx\), and you also have \(x=\dfrac{u+1}{2}\). \[\begin{align*}\int\frac{x}{2x-1}~dx&=\frac{1}{2}\int\frac{\frac{u+1}{2}}{u}~du\\ &=\frac{1}{4}\int\frac{u+1}{u}~du\\ &=\frac{1}{4}\int\left(1+\frac{1}{u}\right)~du\end{align*}\]